# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@ 8.2  Bounding the number of pivot events
Let $z$ be the intersection point of $\ell_{s}$ and $\ell_{t}$. Again refer to Figure˜16 for the following construction. The triangle defined by $s$, $t$ and $z$ does not contain any point of $[s,t]$ in its interior. We transform this triangle by moving $z$ towards $m$ until it hits either $[s,m]$ or $[m,t]$. Without loss of generality, let it hit $[m,t]$. We then continue moving $z$, but away from $t$, instead of towards $m$. The result is a triangle that touches both $(s,m]$ and $[m,t)$, and still does not contain any point of $[s,t]$ in its interior. Let $p_{s}$ and $p_{t}$ be the two points where the triangle touches $(s,m]$ and $[m,t)$. If $p_{s}$ is $m$, then we choose the tangent $h_{s}^{\prime}$ defined by $\ell_{s}^{\prime}$ at $p_{s}$ according to $\pi_{i_{m}}$. If $p_{t}$ is $m$, then we choose the tangent $h_{t}^{\prime}$ defined by $\ell_{t}^{\prime}$ at $p_{t}$ according to $\pi_{i_{m}+1}$. Otherwise, we chose the tangent such that it is colinear with the face of the triangle it is touching. This way, we obtain two tangents that intersect, and their intersection contains the triangle and, in particular, $s$, $t$, $p_{1}$, and $p_{2}$. Further, the line $\ell_{s}^{\prime}$ must intersect $\ell_{1}$ left of $p_{1}$, as otherwise there is a point in $(s,p_{s})$ that is another local minimum of $f$, contradicting the fact that $m$ is the only local minimum of $f$. It follows that, $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ fulfill properties $(i)$, $(ii)$, and $(iii)$, and in particular, a subset of at most three of them as well, concluding the proof. ∎
A consequence of the proof of Lemma˜30 is, that any minimal solution of size at least $4$ contains at least one good guard, strengthening Theorem˜7. We do not make use of this fact. Instead, the consequence we are after is the following.
Corollary 31.
@@
Proof.
This is an immediate consequence of Lemma˜30. ∎
Let us now consider for any vertex $x$ of $P$ the subset $I_{x}\subset\mathcal{U}$ of points $u\in\mathcal{U}$ such that $x$ is on the shortest path $S(u,\textnormal{{next}}(u))$. By Corollary˜31, $u\in L_{x}$ if and only if $x=\min_{\prec^{*}}\{y|y\in I_{u}\}$. Figure˜18 illustrates that the set $L_{x}$ is not necessarily connected, but there is some contiguity for the set $I_{x}$, which we prove in the next lemma.
@@
We conclude that $u^{\prime}$ is left of the second edge of $\pi_{u}$. This implies that both shortest paths $S(u,\textnormal{{next}}(u^{\prime}))$ and $S(u^{\prime},\textnormal{{next}}(u))$ are also left-turning shortest paths that contain $x$. As for any point $\hat{u}\in[u,u^{\prime}]$ we have that $\textnormal{{next}}(\hat{u})\in[\textnormal{{next}}(u),\textnormal{{next}}(u^{\prime})]$, the shortest path $)S(\hat{u},\textnormal{{next}}(\hat{u}$)) cannot properly intersect $S(u,\textnormal{{next}}(u^{\prime}))$ and $S(u^{\prime},\textnormal{{next}}(u))$. Thus, $)S(\hat{u},\textnormal{{next}}(\hat{u}$)) must be left-turning with at least two inner vertices and also have $)x\in S(\hat{u},\textnormal{{next}}(\hat{u}$)). ∎
Corollary 33.
Let $P$ be a simple polygon where no three contiguous guards can guard all of $\partial P$. Then for every vertex $x$ of $P$ there is a contiguous subset $I_{x}^{*}\subset I_{x}\subset\mathcal{U}$ such that $L_{x}\subset I_{x}^{*}$. In particular,
@@
Proof.
This is an immediate consequence of Corollary˜31 and Lemma˜32. ∎
Lemma 34.
@@
Proof.
Consider for every vertex $x$ of $P$ the contiguous subset $I_{x}^{*}\subset I_{x}\subset\mathcal{U}$ as defined in Corollary˜33. These sets define $O(n)$ interior disjoint contiguous intervals in $\mathcal{U}$, where in every interval all points in the interval are contained in the same sets $I_{x}^{*}$. In particular, $u\mapsto\min_{\prec^{*}}\{y|u\in I_{y}^{*}\}$ is constant for each interval. Hence for $L_{x}$, which is the number of connected components of $L_{x}$ as a subset of $\mathcal{U}$, we have:
$\sum_{\text{$x$ vertex of $P$}}\#L_{x}\in O(n).\qed$
Lemma 35.
@@
Proof.
This is an immediate consequence of Lemma˜34, together with the dual nature of $L_{x}$ and $R_{x}$, and the fact that
$\sum_{(i,j,e)\in\mathcal{T}}|\{x|L_{i,j,e,x}\neq\emptyset\}|\leq\left(\sum_{(i,j,e)\in\mathcal{T}}2\right)+\left(\sum_{\text{$x$ vertex of $P$}}\#L_{x}\right)\in O(n).\qed$
Theorem 36.
@@
Proof.
This is an immediate consequence of the computation of all $O(n)$ guards in $\mathcal{D}^{\prime}$, $\mathcal{B}_{\sigma}$ in $O(n\log n)$ time together with Lemma˜25, Lemma˜28, Lemma˜29, and Lemma˜35. ∎
@@ 9  An $O(n\log n)$-time algorithm
Theorem 37.
@@
Proof.
We first check in $O(n\log n)$ time via Corollary˜24, if there is a solution of size at most $3$. Otherwise, via Theorem˜36, we compute the representation of next as a piecewise Möbius transform, consisting of $O(n)$ pieces. Given this representation, there is a $O(n\log k)$ algorithm [aggarwal1989finding], also discussed in [RobsonSpaldingZheng2024_AnalyticArcCover, Remark 10], which can compute the minimum $k$, together with a point $x\in\partial P$, such that $\textnormal{{next}}^{k}(x)$ overtakes $x$, i.e., $x,\textnormal{{next}}^{k}(x)]=\partial P$, or conversely, $\textnormal{{next}}^{k}(x)\geq x+n$. From this $x$, via the representation of next we can compute a solution of size $k$ in time $O(k\log n)$, concluding the proof. ∎
@@ 10  Lower bound
Given two lists $A,B\subset[1,n^{3}]$ of integers, each of size $n$, is $A\cap B=\emptyset$?
Theorem 38.
@@
For ease of arithmetic, we assume that $A$ has $n+1$ elements starting at $A[0]$. For any permutation $\pi$ of $\{1,\ldots,n\}$ we can construct an input $I$ where for all indices $i\in[1,n-1]$, $A[\pi(i)-1]<B[i]<A[\pi(i)]$. For any two permutations $\pi_{1},\pi_{2}$, the resulting certificates must be distinct because for all $i\in[1,n-1]$, the fact that $B[i]\neq A[\pi(i)]$ and $B[i]\neq A[\pi(i)-1]$ cannot be implied by transitivity over comparisons between any $B[j]$ and elements of $A$. It follows that there are $\Theta(n!)$ inputs that all yield different certificates which implies an $\Omega(n\log n)$ comparison-based lower bound. ∎
We now give a construction reducing SetDisjointness to computing the minimum number of guards for a given polygon. For this we construct three gadgets, that we place on the boundary of a big polygon. The big polygon is a right isosceles triangle, whose base has length $n^{3}$. For the following construction refer to Figure˜19.
@@
Since $A$ is sorted, we can perform the first part of the construction in linear time, describing the polygon in a clock-wise order. The remaining part of the construction can also be completed in a clock-wise order: the disjoiners do not rely on the input, and the laser gadgets can be constructed by taking the line from $(b,\min(b,n^{3}-b))$ through four points on the $x$-axis, and intersecting with some well-chosen rectangle below the $x$-axis.
Lemma 39.
No contiguous guard can guard two points $u$ and $v$ between which a disjoiner gadget lies. I.e., no contiguous guard can guard a chain $[u,v]$ where the open chain $(u,v)$ that contains a disjoiner gadget. Also, no contiguous guard can guard two disjoiner gadgets at once.
Proof.
@@
As a contiguous guard that sees an interior disjoiner gadget has to be placed inside the rectangle defining the gadget, and no guard can be placed in the rectangle of two different disjoiner gadget, no two disjoiner gadgets can be guard by the same contiguous guard. ∎
Lemma 40.
@@
Now suppose, a solution of size $2n+2$ exists. By Lemma˜39, at least $n+1$ of those must guard the $n+1$ areas of the polygon that are separated by disjoiner gadgets. That leaves $n+1$ unguarded disjoiner gadgets, which each require their own guard. Hence each area that is separated by a disjoiner gadget has to be guarded by a single guard. In particular, every laser gadget has to be guarded by a single guard. But this is only possible, if $(b,\min(b,n^{3}-b))$ is in the polygon for every $b\in B$. This in turn is only possible, if $B\cap A=\emptyset$, concluding the proof. ∎
@@
Proof.
This is an immediate consequence of Theorem˜38 and Lemma˜40, together with the fact that the construction from Lemma˜40 takes $O(n)$ time, and the size of the coordinates of the vertices is polynomial in the input size. ∎
Remark 42.
@@ 11  Conclusion
We presented matching upper and lower bounds for the Contiguous Art Gallery problem.
A natural direction for further research is to investigate whether our algorithmic ideas extend to polygons with holes. For polygons with holes, there are two natural variants: either one must guard the boundaries of the holes in addition to the outer boundary, or only the outer boundary is required to be guarded. In both cases, at least the outer boundary must be covered. For our approach, however, several crucial components break down when holes are allowed. Not only do key data structures, such as the shortest-path structure, no longer behave as required, but certain structural properties also fail to generalise. In particular, Theorem˜7 shows that for simple polygons any solution can be modified so that it contains a good guard. Given the $O(n)$ reduced good dominators, implementing next yields the $O(kn\log n)$ algorithm. In polygons with holes, by contrast, there exist configurations in which no such modification is possible (see Figure˜21). The essential insight behind the proof of Theorem˜7 is that every bad guard is bounded by reflex vertices of the outer boundary, which themselves must be guarded. If instead the bounding vertices occur on holes, we can no longer force the presence of a good guard. It is therefore likely that new structural insights are required to obtain near-linear-time algorithms for polygons with holes, if near-linear time is achievable at all.
@@
Figure 20: Illustration of a set of guards guarding the outer boundary of a polygon $P$ with a hole. The solution cannot readily be transformed via Theorem˜7 to have at least one good guard.

### v4

@@ 8.2 Bounding the number of pivot events
Let $z$ be the intersection point of $\ell_{s}$ and $\ell_{t}$. Again refer to Figure˜16 for the following construction. The triangle defined by $s$, $t$ and $z$ does not contain any point of $[s,t]$ in its interior. We transform this triangle by moving $z$ towards $m$ until it hits either $[s,m]$ or $[m,t]$. Without loss of generality, let it hit $[m,t]$. We then continue moving $z$, but away from $t$, instead of towards $m$. The result is a triangle that touches both $(s,m]$ and $[m,t)$, and still does not contain any point of $[s,t]$ in its interior. Let $p_{s}$ and $p_{t}$ be the two points where the triangle touches $(s,m]$ and $[m,t)$. If $p_{s}$ is $m$, then we choose the tangent $h_{s}^{\prime}$ defined by $\ell_{s}^{\prime}$ at $p_{s}$ according to $\pi_{i_{m}}$. If $p_{t}$ is $m$, then we choose the tangent $h_{t}^{\prime}$ defined by $\ell_{t}^{\prime}$ at $p_{t}$ according to $\pi_{i_{m}+1}$. Otherwise, we chose the tangent such that it is colinear with the face of the triangle it is touching. This way, we obtain two tangents that intersect, and their intersection contains the triangle and, in particular, $s$, $t$, $p_{1}$, and $p_{2}$. Further, the line $\ell_{s}^{\prime}$ must intersect $\ell_{1}$ left of $p_{1}$, as otherwise there is a point in $(s,p_{s})$ that is another local minimum of $f$, contradicting the fact that $m$ is the only local minimum of $f$. It follows that, $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ fulfill properties $(i)$, $(ii)$, and $(iii)$, and in particular, a subset of at most three of them as well, concluding the proof.
A consequence of the proof of Lemma˜8.10 is, that any minimal solution of size at least $4$ contains at least one good guard, strengthening Theorem˜4.17. We do not make use of this fact. Instead, the consequence we are after is the following.
Corollary 8.12.
@@
Proof 8.13.
This is an immediate consequence of Lemma˜8.10.
Let us now consider for any vertex $x$ of $P$ the subset $I_{x}\subseteq\mathcal{U}$ of points $u\in\mathcal{U}$ such that $x$ is on the shortest path $S(u,\textnormal{{next}}(u))$. By Corollary˜8.12, $u\in L_{x}$ if and only if $x=\min_{\prec^{*}}\{y|y\in I_{u}\}$. Figure˜18 illustrates that the set $L_{x}$ is not necessarily connected, but there is some contiguity for the set $I_{x}$, which we prove in the next lemma.
@@
We conclude that $u^{\prime}$ is left of the second edge of $\pi_{u}$. This implies that both shortest paths $S(u,\textnormal{{next}}(u^{\prime}))$ and $S(u^{\prime},\textnormal{{next}}(u))$ are also left-turning shortest paths that contain $x$. As for any point $\hat{u}\in[u,u^{\prime}]$ we have that $\textnormal{{next}}(\hat{u})\in[\textnormal{{next}}(u),\textnormal{{next}}(u^{\prime})]$, the shortest path $)S(\hat{u},\textnormal{{next}}(\hat{u}$)) cannot properly intersect $S(u,\textnormal{{next}}(u^{\prime}))$ and $S(u^{\prime},\textnormal{{next}}(u))$. Thus, $)S(\hat{u},\textnormal{{next}}(\hat{u}$)) must be left-turning with at least two inner vertices and also have $)x\in S(\hat{u},\textnormal{{next}}(\hat{u}$)).
Corollary 8.16.
Let $P$ be a simple polygon where no three contiguous guards can guard all of $\partial P$. Then for every vertex $x$ of $P$ there is a contiguous subset $I_{x}^{*}\subseteq I_{x}\subseteq\mathcal{U}$ such that $L_{x}\subseteq I_{x}^{*}$. In particular,
@@
Proof 8.17.
This is an immediate consequence of Corollary˜8.12 and Lemma˜8.14.
Lemma 8.18.
@@
Proof 8.19.
Consider for every vertex $x$ of $P$ the contiguous subset $I_{x}^{*}\subseteq I_{x}\subseteq\mathcal{U}$ as defined in Corollary˜8.16. These sets define $O(n)$ interior disjoint contiguous intervals in $\mathcal{U}$, where in every interval all points in the interval are contained in the same sets $I_{x}^{*}$. In particular, $u\mapsto\min_{\prec^{*}}\{y|u\in I_{y}^{*}\}$ is constant for each interval. Hence for $L_{x}$, which is the number of connected components of $L_{x}$ as a subset of $\mathcal{U}$, we have:
$\sum_{\text{$x$ vertex of $P$}}\#L_{x}\in O(n).$
Lemma 8.20.
@@
Proof 8.21.
This is an immediate consequence of Lemma˜8.18, together with the dual nature of $L_{x}$ and $R_{x}$, and the fact that
$\sum_{(i,j,e)\in\mathcal{T}}|\{x|L_{i,j,e,x}\neq\emptyset\}|\leq\left(\sum_{(i,j,e)\in\mathcal{T}}2\right)+\left(\sum_{\text{$x$ vertex of $P$}}\#L_{x}\right)\in O(n).$
Theorem 8.22.
@@
Proof 8.23.
This is an immediate consequence of the computation of all $O(n)$ guards in $\mathcal{D}^{\prime}$, $\mathcal{B}_{\sigma}$ in $O(n\log n)$ time together with Lemma˜8.1, Lemma˜8.6, Lemma˜8.8, and Lemma˜8.20.
@@ 9 An $O(n\log n)$-time algorithm
Given the presented results, we can finally present our main result. After $O(n\log n)$ time we can either output a solution of size at most $3$, or have a representation of next as a piecewise Möbius transform, consisting of $O(n)$ pieces. Note that given this data, we can get a $(+1)$-approximation $k^{\prime}$ of $k$ in $O(n)$ time by repeatedly applying $\textnormal{{next}}(\cdot)$ to an arbitrary $x\in\partial P$ [RobsonSpaldingZheng2024_AnalyticArcCover, Lemma 7] via a linear sweep of the representation of next. We finally use the techniques of [aggarwal1989finding, Section 4], which in $O(n\log k)$ checks if there is any point $x\in\partial P$ such that $\textnormal{{next}}^{k^{\prime}-1}(x)\geq x+n$.
Theorem 9.1.
@@
Proof 9.2.
[…]
@@ 10 Lower bound
Given two lists $A,B\subseteq[1,n^{3}]$ of integers, each of size $n$, is $A\cap B=\emptyset$?
Theorem 10.1.
@@
For ease of arithmetic, we assume that $A$ has $n+1$ elements starting at $A[0]$. For any permutation $\pi$ of $\{1,\ldots,n\}$ we can construct an input $I$ where for all indices $i\in[1,n-1]$, $A[\pi(i)-1]<B[i]<A[\pi(i)]$. For any two permutations $\pi_{1},\pi_{2}$, the resulting certificates must be distinct because for all $i\in[1,n-1]$, the fact that $B[i]\neq A[\pi(i)]$ and $B[i]\neq A[\pi(i)-1]$ cannot be implied by transitivity over comparisons between any $B[j]$ and elements of $A$. It follows that there are $\Theta(n!)$ inputs that all yield different certificates which implies an $\Omega(n\log n)$ comparison-based lower bound.
We now give a construction reducing SetDisjointness where $A$ and $B$ are integer sets and $A$ is sorted to computing the minimum number of guards for a given polygon. For this we construct three gadgets, that we place on the boundary of a big polygon. The big polygon is a right isosceles triangle, whose base has length $n^{3}$. For the following construction refer to Figure˜19.
@@
Since $A$ is sorted, we can perform the first part of the construction in linear time, describing the polygon in a clock-wise order. The remaining part of the construction can also be completed in a clock-wise order in linear time: the disjoiners do not rely on the input, and the laser gadgets can be constructed by taking the line from $(b,\min(b,n^{3}-b))$ through four points on the $x$-axis, and intersecting with some well-chosen rectangle below the $x$-axis. The laser gadgets do not rely on $B$ being sorted and can be constructed in the given order of $B$ along the polygon boundary (refer to the crossing bold red construction lines in Figure˜19).
Lemma 10.3.
No contiguous guard can guard two points $u$ and $v$ between which a disjoiner gadget lies. I.e., no contiguous guard can guard a chain $[u,v]$ where the open chain $(u,v)$ contains a disjoiner gadget. Also, no contiguous guard can guard two disjoiner gadgets at once.
Proof 10.4.
@@
As a contiguous guard that sees an interior disjoiner gadget has to be placed inside the rectangle defining the gadget, and no guard can be placed in the rectangle of two different disjoiner gadget, no two disjoiner gadgets can be guard by the same contiguous guard.
Lemma 10.5.
@@
Now suppose, a solution of size $2n+2$ exists. By Lemma˜10.3, at least $n+1$ of those must guard the $n+1$ areas of the polygon that are separated by disjoiner gadgets. That leaves $n+1$ unguarded disjoiner gadgets, which each require their own guard. Hence each area that is separated by a disjoiner gadget has to be guarded by a single guard. In particular, every laser gadget has to be guarded by a single guard. But this is only possible, if $(b,\min(b,n^{3}-b))$ is in the polygon for every $b\in B$. This in turn is only possible, if $B\cap A=\emptyset$, concluding the proof.
@@
Proof 10.8.
This is an immediate consequence of Theorem˜10.1 and Lemma˜10.5, together with the fact that the construction from Lemma˜10.5 takes $O(n)$ time, and the size of the coordinates of the vertices is polynomial in the input size.
Remark 10.9.
@@ 11 Conclusion
We presented matching upper and lower bounds for the Contiguous Art Gallery problem. A natural direction for further research is to investigate whether our algorithmic ideas extend to polygons with holes. For polygons with holes, there are two natural variants: either one must guard the boundaries of the holes in addition to the outer boundary, or only the outer boundary is required to be guarded. In both cases, at least the outer boundary must be covered. For our approach, however, several crucial components break down when holes are allowed. Not only do key data structures, such as the shortest-path structure, no longer behave as required, but certain structural properties also fail to generalise. In particular, Theorem˜4.17 shows that for simple polygons any solution can be modified so that it contains a good guard. Given the $O(n)$ reduced good dominators, implementing next yields the $O(kn\log n)$ algorithm. In polygons with holes, by contrast, there exist configurations in which no such modification is possible (see Figure˜21). The essential insight behind the proof of Theorem˜4.17 is that every bad guard is bounded by reflex vertices of the outer boundary, which themselves must be guarded. If instead the bounding vertices occur on holes, we can no longer force the presence of a good guard. It is therefore likely that new structural insights are required to obtain near-linear-time algorithms for polygons with holes, if near-linear time is achievable at all.
@@
Figure 20: Illustration of a set of guards guarding the outer boundary of a polygon $P$ with a hole. The solution cannot readily be transformed via Theorem˜4.17 to have at least one good guard.